# Adaptive Rank Allocation for Federated Parameter- Efficient Fine-Tuning of Language Models

### v2

@@ VI-B Training Time on Advanced Edge Devices
We evaluate the training time performance of FedARA and other baselines on various edge devices. On Raspberry Pi 5 (RPi 5), local training times per batch (batch size = 4) are 1.00 second for the DistilBERT model and 2.01 seconds for the BERT model. AGX Orin and Orin Nano exhibit significantly faster training times under the DistilBERT model, achieving 6.67$\times$ and 5.56$\times$ speedups over RPi 5, respectively. For the BERT model, these differences increases to 8.74$\times$ and 6.70$\times$.
@@
Figure 9 compares the total training time for the DistilBERT model across these devices. On RPi 5, FedARA reduces the average total training time to 68 minutes, representing a 27.65% and 24.38% reduction compared to FedLoRA and FedLoRA-B-dr, respectively. On Orin Nano and AGX Orin, FedARA achieves training times of 25 and 23 minutes, corresponding to reductions of 46.57% and 48.90% relative to FedLoRA, and 8.99% and 12.20% compared to FedLoRA-B. The results indicate that the limited computational capability of edge devices may diminish the benefits of reduced communication overhead. These findings align with preliminary observations (Figure 2d) that high-end devices are mainly constrained by communication, while computation remains a bottleneck for lower-end devices.
FedARA achieves notable training time improvements through dynamic rank allocation and rank-based module pruning. Dynamic rank allocation significantly reduces communication overhead by minimizing transmitted data, while rank-based module pruning further optimizes local training by structurally removing inactive SVD modules, thereby decreasing computation time and resource consumption on local devices.
@@ VI-C Ablation Experiments
Ablation study of rank-based module pruning. Rank-based module pruning does not impact performance, as it preserves the same parameters for communication and global aggregation. We evaluate its effect on local training time and GPU usage on Semeval and News Category datasets. As illustrated in Figure 12 and 13, while FedARA initially exhibits a slight increase in local training time, it significantly reduces over time. Compared to FedLoRA, FedARA lowers average local training time by 10.81% and reduces GPU memory consumption by 31.67% per round, demonstrating its efficiency in minimizing computational and storage overhead.
@@
Figure 14: Impact of different target rank and arbitration threshold settings on validation loss and communication overhead. The upper sub-figure presents results on the 20News dataset, while the lower sub-figure corresponds to AG News (Model: DistilBERT).
VI-D Extended Experiments
To further evaluate the impact of the pre-set target average rank ($T_{r}$) and arbitration threshold ($T_{h}$) on FedARA, we conduct extended experiments where the initial rank is set to 12 and a default arbitration threshold of 0.5. As shown in the upper part of Figure 14, lower target ranks reduce communication overhead but compromise validation loss, while higher ranks enhance performance at the expense of increased overhead. To balance performance and communication efficiency, we set the predefined target average rank ($T_{r}$) to one-fourth of the initial rank in this study.
Additionally, the lower part of Figure 14 illustrates that as the arbitration threshold ($T_{h}$) increases, the validation loss exhibits a slight upward trend, while communication overhead marginally decreases. This occurs because higher thresholds impose stricter arbitration, causing more ranks to be classified as less important. Notably, FedARA remains insensitive to the arbitration threshold, suggesting that local rank masks from different clients converge toward similarity during training. Based on these findings, we adopt 0.5 as the default arbitration threshold ($T_{h}$) in this work.
@@ VI-E Local Resource Utilization
Memory Footprint. Although SVD adaptation introduces an additional diagonal matrix, the peak memory overhead remains negligible due to the significantly smaller rank dimension $r$ compared to the hidden layer size. As shown in Figure 15, integrating SVD modules and dynamic rank allocation in FedARA results in only a marginal increase in peak memory footprint compared to FedLoRA, with 2.42% for DistilBERT and 4.40% for BERT on Orin Nano. More importantly, FedARA progressively reduces memory consumption per round during federated fine-tuning. These results confirm that FedARA is well-suited for memory-constrained devices.
Energy Consumption. Figure 16 presents the energy consumption of different methods on Orin Nano (15 Watts power) over 100 rounds with 10 clients per round. FedARA achieves the lowest average energy consumption, with approximately $2.3\times 10^{5}$ Joules for the DistilBERT model and $4.4\times 10^{5}$ Joules for the BERT model, marking a 46.95% reduction compared to FedLoRA and about 9% reduction compared to FedLoRA-B. This highlights that FedARA can significantly reduce energy consumption to prolong device runtime and decrease operational costs for sustainable deployment in resource-constrained devices.
@@ VII Related Work
FedARA builds upon advancements in Parameter-Efficient Fine-Tuning (PEFT) and its extension to Federated Parameter-Efficient Fine-Tuning (FedPEFT). Therefore, this section investigates key developments in PEFT and FedPEFT and explores related efforts to address two key challenges in FedPEFT: mitigating non-IID data effects and enabling dynamic parameter configurations.
PEFT. PEFT enables effective and efficient adaptation to diverse downstream tasks by fine-tuning only a small subset of parameters, significantly reducing resource consumption compared to full fine-tuning. Based on the insertion positions and integration mechanisms of PEFT modules, PEFT methods are categorized as follows: 1) Adapter-based methods [11, 12], which sequentially insert small adapter modules after the output linear layers of transformers. 2) Prompt-based methods [13, 14], which place trainable parameters into the input prompt. 3) LoRA-based methods [15], which integrate low-rank modules in parallel with the linear layers of transformers. The proposed FedARA adopts LoRA-based PEFT modules for local training, enabling seamless integration into PLMs without introducing additional inference latency.
@@
FedPEFT. FedPEFT extends PEFT methods to the federated learning framework, leveraging PEFT’s efficiency to alleviate the substantial communication overhead associated with FL on PLMs. Additionally, FedPEFT offers promising solutions to unique FL challenges. FedIT [16] enables privacy-preserving acquisition of high-quality instruction data for generative tasks. FLoRA [17] introduces a noise-free aggregation method tailored for heterogeneous FL settings. FlexLoRA [19] presents an efficient FL framework designed for multi-task learning and heterogeneous resource environments. Moreover, established studies have also advanced the system implementation of FedPEFT [25, 26]. The proposed FedARA aims to mitigate the performance degradation of FedPEFT methods under non-IID settings and enhance communication efficiency by addressing the inefficiencies of fixed parameter configurations.
Non-IID Data. Non-IID data in FL, where clients have distinct distributions, complicates model aggregation and hinders global generalization. Traditional strategies such as client selection, optimization, and aggregation have proven effective in mitigating these effects [33, 34, 35, 36, 37, 38, 39, 40]. Recent studies have also focused on addressing the challenges of non-IID data in pre-trained models within the PEFT framework. SLoRA [20] adopts a two-stage training approach: it first applies sparse fine-tuning to establish a better initialization, and then it utilizes LoRA adaptation based on the results of sparse fine-tuning. FeDeRA [21] leverages SVD decomposition to optimize LoRA module initialization. These methods retain the original LoRA structure, relying solely on improved parameter initialization. Beyond initialization strategies, recent studies explore asymmetrical training that only updates matrix $B$ and frozen matrix $A$ in LoRA to mitigate non-IID effects [28, 29, 24], leveraging the distinct functions of LoRA’s two matrices. However, these methods require a delicate trade-off between expressiveness and the number of trainable parameters. The proposed FedARA directly changes the structure and initialization with truncated SVD adaptation, improving its ability to represent similar features across different clients.
Dynamic Parameter Configurations. Fixed parameter configurations lead to suboptimal performance and communication inefficiency, making dynamic parameter allocation a superior alternative. In centralized training, AdaLoRA [32] introduces a sensitivity-based dynamic rank allocation method for LoRA. In the context of FedPEFT, SA-FedLoRA [23] adopts a two-stage simulated annealing-based approach: the first stage performs several rounds of full fine-tuning with a parameter regularization penalty to prevent model drift, and the second stage adaptively adjusts the ranks of LoRA using simulated annealing. However, the high computational cost of full fine-tuning makes this method impractical for resource-constrained devices. Similarly, AdaFL [25] dynamically expands parameter sizes, with the growth direction determined by sideline trials conducted on other clients. However, this adapter-based approach introduces additional inference latency, and sideline trials further increase communication costs for extra clients. To address these limitations, FedARA replaces full fine-tuning with truncated SVD adaptation, making it feasible for resource-constrained devices. By leveraging rank masks, FedARA progressively reduces communication overhead and eliminates the need for additional clients, offering a more efficient and scalable solution.
@@ VIII Conclusion
FedARA is a novel framework for Federated Parameter-Efficient Fine-Tuning (FedPEFT) of language models. It enhances performance under non-IID data and improves communication efficiency by integrating adaptive rank allocation with truncated SVD modules. Additionally, FedARA minimizes local resource usage per round in FL. Extensive experiments demonstrate its superior performance and efficient communication compared to baselines, establishing FedARA as a promising solution for FedPEFT. For future work, we plan to extend FedARA to a broader range of domains and tasks with larger models while exploring its integration with quantization techniques to further facilitate the adaptability of pre-trained models for diverse downstream tasks on resource-constrained devices.

### v3

@@ VI-B Training Time on Advanced Edge Devices
We evaluate the training efficiency of FedARA and baselines on various edge devices. On Raspberry Pi 5 (RPi 5), local training time per batch (batch size = 4) is 1.00 s for DistilBERT and 2.01 s for BERT. AGX Orin and Orin Nano achieve significantly faster training under DistilBERT, with 6.67$\times$ and 5.56$\times$ speedups over RPi 5, respectively; for BERT, the speedups increase to 8.74$\times$ and 6.70$\times$.
@@
Figure 10 compares total training time for the DistilBERT model across devices. On RPi 5, FedARA reduces training time to 68 minutes, representing 27.65% and 24.38% reductions over FedLoRA and FFA-LoRA-dr, respectively. On Orin Nano and AGX Orin, FedARA further lowers training time to 25 and 23 minutes, achieving up to 48.90% reduction over FedLoRA and over 12% compared to FFA-LoRA. These results suggest that limited compute on edge devices may diminish the benefits of reduced communication overhead. This observation aligns with Figure 2d, where high-end devices are communication-bound, while low-end devices remain compute-bound. We exclude comparisons with SLoRA and FeDeRA due to their noticeably higher training and computational costs compared to FedLoRA. Specifically, SLoRA incurs overhead comparable to full fine-tuning when using optimizer masking approach [18], while FeDeRA introduces additional model-size-dependent costs through SVD.
@@ VI-C Ablation Experiments
Ablation study of rank-based module pruning. Rank-based module pruning does not impact performance, as it preserves the same parameters for communication and global aggregation. We evaluate its effect on local training time and GPU usage on Semeval and News Category datasets. As illustrated in Figure 13 and 14, while FedARA initially exhibits a slight increase in local training time, it significantly reduces over time. Compared to FedLoRA, FedARA lowers average local training time by 10.81% and reduces GPU memory consumption by 31.67% per round, demonstrating its efficiency in minimizing computational and storage overhead.
VI-D Extended Experiments
To assess the impact of the target average rank ($T_{r}$) and arbitration threshold ($T_{h}$), we conduct extended experiments with an initial rank of 12 and a default threshold of 0.5. As shown in the upper part of Figure 15, lower target ranks reduce communication overhead but compromise validation loss, while higher ranks enhance performance at the expense of increased overhead. To balance performance and communication efficiency, we set the predefined target average rank ($T_{r}$) to one-fourth of the initial rank in this study.
@@ VI-D Extended Experiments
Figure 15: Impact of target rank and arbitration threshold on validation loss and communication overhead. The upper subfigure shows results on 20News, and the lower on AG News using the DistilBERT model.
Additionally, the lower part of Figure 15 shows that increasing the arbitration threshold ($T_{h}$) slightly increases validation loss while marginally reducing communication overhead. This is because a higher $T_{h}$ enforces stricter arbitration, leading to more ranks being deemed unimportant. Notably, FedARA is relatively insensitive to $T_{h}$, indicating that local rank masks tend to converge across clients. Based on these observations, we adopt 0.5 as the default value for $T_{h}$ in this study.
@@ VI-E Local Resource Utilization
Memory Footprint. Although SVD adaptation introduces an additional diagonal matrix, the memory overhead remains negligible due to the small rank $r$ relative to the hidden layer size. As shown in Figure 16, integrating SVD modules and dynamic rank allocation in FedARA leads to only a slight increase in peak memory usage compared to FedLoRA—2.42% for DistilBERT and 4.40% for BERT on Orin Nano. Moreover, FedARA progressively reduces memory usage per round during federated fine-tuning, confirming its suitability for memory-constrained devices.
Energy Consumption. Figure 17 shows the energy consumption on Orin Nano (15W) over 100 rounds with 10 clients per round. FedARA achieves the lowest energy usage—approximately $2.3\times 10^{5}$ J for DistilBERT and $4.4\times 10^{5}$ J for BERT—representing a 46.95% reduction compared to FedLoRA and around 9% compared to FFA-LoRA. These results demonstrate FedARA’s ability to reduce energy consumption, extending device runtime and lowering deployment costs for sustainable use in resource-constrained environments.
@@ VII Related Work
FedARA builds on recent advances in Parameter-Efficient Fine-Tuning (PEFT) and its extension to the federated PEFT (FedPEFT). This section reviews key developments in both areas and examines related work addressing two main challenges in FedPEFT: mitigating non-IID effects and supporting dynamic parameter configurations.
PEFT. PEFT enables efficient adaptation to downstream tasks by fine-tuning only a small subset of parameters, significantly reducing resource consumption compared to full fine-tuning. Based on insertion positions and integration mechanisms, PEFT methods can be categorized into: 1) adapter-based methods [11, 12], which insert small modules after transformer linear layers; 2) prompt-based methods [13, 14], which embed trainable parameters into input prompts; and 3) LoRA-based methods [15], which inject low-rank matrices in parallel with transformer linear layers. FedARA adopts LoRA-based modules for local training, allowing seamless integration into PLMs without additional inference latency.
FedPEFT. FedPEFT extends PEFT to federated learning, leveraging its efficiency to reduce the substantial communication overhead of applying PEFT to PLMs. FedIT [16] facilitates privacy-preserving acquisition of high-quality instruction data for generative tasks. FLoRA [17] proposes a noise-free aggregation strategy tailored to heterogeneous FL, while FlexLoRA [18] presents an efficient framework for multi-task FL under resource heterogeneity. Other efforts focus on system-level optimization of FedPEFT [23, 24] and open-source frameworks for LLMs and heterogeneous edge environments [44, 45]. These works offer valuable context for real-world federated LLM deployment and complement our focus on efficient and adaptive federated fine-tuning under non-IID conditions.
Non-IID Data. Non-IID data in FL hinders model aggregation and global generalization. Traditional methods—such as client selection, optimization, and aggregation—have been effective in mitigating these effects [29, 30, 31, 32, 33, 34, 35, 36]. Recent work extends this line to pre-trained models within the FedPEFT framework. SLoRA [19] adopts a two-stage strategy: sparse fine-tuning for initialization, followed by LoRA adaptation. FeDeRA [20] enhances LoRA initialization via SVD. Beyond initialization, recent studies explore asymmetric training by updating only matrix $B$ while freezing matrix $A$ to reduce non-IID impact [43, 26, 22]. FedARA further addresses non-IID challenges by modifying both the structure and initialization via truncated SVD adaptation, effectively slowing client drift.
@@
Dynamic Parameter Configurations. Fixed rank configurations lead to dual losses of suboptimal performance and communication inefficiency. In centralized training, AdaLoRA [28] introduces a sensitivity-based dynamic rank allocation method. In the context of FedPEFT, AdaFL [23] expands parameters via sideline trials on other clients, but introduces inference latency and additional communication costs. SA-FedLoRA [21] employs a two-stage simulated annealing approach: full fine-tuning with regularization to control model drift, followed by adaptive LoRA rank adjustment. However, the high cost of full fine-tuning limits its practicality on resource-constrained devices. FedARA replaces full fine-tuning with truncated SVD adaptation for resource efficiency, and leverages rank masks to reduce communication without extra clients, enabling a more scalable solution.
@@ VIII Conclusion
FedARA is a novel framework for Federated Parameter-Efficient Fine-Tuning (FedPEFT) of language models. It improves performance under non-IID data and improves communication efficiency by integrating adaptive rank allocation with truncated SVD adaptation, while minimizing per-round local resource usage. Extensive experiments validate its superiority over existing baselines. In future work, we aim to extend FedARA to larger models and broader tasks, and explore integration with quantization to further improve adaptability on resource-constrained devices.